OpenStudy (itsonlycdeee):

Find the derivative of: f(x)=[ x^(1/2)+1]/[x^2 + 1]

OpenStudy (shamil98):

use the power rule, for 1/x^2 use this: and then the power rule \[\large \frac{ 1 }{ x^n } = x^{-n}\] the derivative of a constant is zero.

OpenStudy (itsonlycdeee):

Thank you, so you're splitting them up? & for some odd reason, what you wrote isn't showing up :/

OpenStudy (shamil98):

oh, I think it's because OpenStudy isn't showing LaTeX for everyone.

OpenStudy (shamil98):

f(x) = x^(1/2) + x^(-2) + 1 power rule is f(x) = x^n f'(x) = nx^(n-1)

OpenStudy (itsonlycdeee):

Ohhhh, I get what you did. You're making it all one line instead of using the quotient rule right?

OpenStudy (shamil98):

all you do is differentiate each value with respect to x . so f(x) = x^(1/2) + x^(-2) + 1 f'(x) = d/dx (x^(1/2) + d/dx (x^(-2)) + d/dx (1) yeah, i find it easier than having to use the quotient rule.

OpenStudy (itsonlycdeee):

Okay thank you

OpenStudy (shamil98):

x^-n = 1/x^n x^-2 = 1/x^2 np